3. A baseball team has 9 players, all of whom have equal skill in all positions. The team needs 3 outfielders. Assuming order is not important, how many possible combinations of outfielders are there? Put your formula/answer in cell A3.

4. The results of a recent survey indicate that 68% of households in the United States have internet access. If you randomly select 5 households, what is the probability that exactly 3 households will have internet access? Put your answer in cell A4. What is the probability that fewer than 3 households have internet access? Put your answer in cell A5. What is the probability at least three of them have internet access? Put your answer in cell A6. (False returns “exactly” amount”, whereas “true” returns the cumulative amount.)

5. Find the area under the standard normal curve to the left of z = -1.01. Put your answer in cell A7.
: 3. A baseball team has 9 players, all of whom have equal skill in all positions. The team needs 3 outfielders. Assuming order is not important, how many possible combinations of outfielders are there? Put your formula/answer in cell A3.

4. The results of a recent survey indicate that 68% of households in the United States have internet access. If you randomly select 5 households, what is the probability that exactly 3 households will have internet access? Put your answer in cell A4. What is the probability that fewer than 3 households have internet access? Put your answer in cell A5. What is the probability at least three of them have internet access? Put your answer in cell A6. (False returns “exactly” amount”, whereas “true” returns the cumulative amount.)

5. Find the area under the standard normal curve to the left of z = -1.01. Put your answer in cell A7.

Answer by stanbon(18792)

3. A baseball team has 9 players, all of whom have equal skill in all positions. The team needs 3 outfielders. Assuming order is not important, how many possible combinations of outfielders are there? Put your formula/answer in cell A3.
Ans: 9C3 = 9*8*7/1*2*3 = 3*4*7 = 84
----------------------
Comment: I am using a TI calculator to get the following results.
You may get slightly different results using EXCEL.
----------------------
4. The results of a recent survey indicate that 68% of households in the United States have internet access.
If you randomly select 5 households, what is the probability that exactly 3 households will have internet access?
Ans: 5C3(0.68)^3(0.32)^2 = 0.32197...
Put your answer in cell A4.
------------------------
What is the probability that fewer than 3 households have internet access? Put your answer in cell A5.
Ans: binomcdf(5,0.68,2) = 0.190526...
----------------------------
What is the probability at least three of them have internet access? Put your answer in cell A6. (False returns “exactly” amount”, whereas “true” returns the cumulative amount.)
Ans: 1 - binomcdf(5,0.68,2) = 0.80947....
--------------------------------
5. Find the area under the standard normal curve to the left of z = -1.01. Put your answer in cell A7.
Ans: P(-10 < z < -1,01) = normalcdf(-10,-1.01) = 0.15624...
======================
Cheers,
Stan H.